# Perverse microsheaves

### v2

@@ 4.1. Sheaves on manifolds
Let $M$ be a real manifold. Fix a symmetric monoidal stable presentable category , $C$. The reader will not lose much of the point of the paper taking throughout $C$ to be the derived category of dg modules over some commutative ring $R$. We write $sh(M)$ for the (stable) category of sheaves on $M$ with values in $C$.
@@ 4.1.1. Microsupport
The microsupport is easily seen to be conical and satisfy $ss(Cone(F\to G))\subset ss(F)\cup ss(G)$. A deep result of [15, Thm. 6.5.4] is that the microsupport is coisotropic (also called involutive; see [15, Def. 6.5.1]). For a conic subset $K\subset T^{*}M$, we write $sh_{K}(M)$ for the full subcategory on objects microsupported in $K$. For any subset $\Lambda\subset S^{*}M$, we write $sh_{\Lambda}(M):=sh_{\mathbb{R}_{+}\Lambda\cup 0_{M}}(M)$, with $0_{M}$ the zero section of the cotangent bundle.
@@ 4.2. Microsheaves on cotangent bundles
Let $\mu sh_{T^{*}M}$ be the sheaf of categories on $T^{*}M$ defined by sheafifying the presheaf $sh_{T^{*}M}^{pre}(U)$ in (9). Similarly, let $\mu sh_{S^{*}M}$ be the presheaf of categories on $S^{*}M$ obtained by sheafifying $\mu sh^{pre}_{S^{*}M}$.
@@
The sheaf $\mu sh_{T^{*}M}$ is conic, i.e. equivariant for the $\mathbb{R}^{+}$ scaling action. In particular, $\mu sh_{T^{*}M}|_{T^{\circ}M}$ is locally constant in the radial direction. This being a contractible $\mathbb{R}^{+}$, we may define a sheaf of categories $\mu sh_{S^{*}M}$ on $S^{*}M$ equivalently by pushforward or pullback along an arbitrary section of $T^{\circ}M\to S^{*}M$.
@@
Let $M$ be a manifold and let $N\subset M$ be a submanifold. Let $L=T^{*}_{N}M$ and fix a point $p\in L$. Then there is an equivalence of categories:
(10) $\omega_{p}:C$ $\xrightarrow{\sim}(\mu sh_{L,\phi})_{p}$
@@
Remark 4.7.
Obviously $\omega_{p}$ depends only on the germ of $M$ near the image of $p$; for this remark we regard $M,N$ as germs. Consider another triple $M^{\prime},N^{\prime},L^{\prime}=T^{*}_{N^{\prime}}M^{\prime},p^{\prime}\in L^{\prime}$. If $\dim M=\dim M^{\prime}$ and $\dim N=\dim N^{\prime}$, then there is a diffeomorphism $(M,N)\to(M^{\prime},N^{\prime})$ whose canonical lift to the cotangent bundle sends $p\to p^{\prime}$. (Indeed, note that $p$ may be moved arbitrarily in $L\setminus 0$ by diffeomorphisms of $M$ fixing $N$.) Obviously the construction of $\omega_{p}$ is functorial in such diffeomorphisms.
In any sheaf of categories $\mathcal{X}$ on a topological space $T$, given $F,G\in\mathcal{X}(T)$, the assignment $U\mapsto\operatorname{Hom}_{\mathcal{X}(U)}(F|_{U},G|_{U})$ is a sheaf on $T$; let us denote it as $\mathcal{H}om_{\mathcal{X}}(F,G)$. Per [15, Thm. 6.1.2]: [^1] More precisely, [15] show there is a morphism of this kind for $\mu sh^{pre}$ which is an isomorphism at stalks; the stated result follows upon sheafification. See [22] for some detailed discussions about the sheafification of $\mu sh^{pre}$.
(11) $\mathcal{H}om_{\mu sh}(F,G)=\mu hom(F,G)$
where $\mu hom$ is a functor explicitly defined in terms of standard sheaf theoretic operations, namely the Fourier-Sato transformation along the diagonal of the external Hom:
$\mu hom(F,G):=\mu_{\Delta}\operatorname{Hom}_{M\times M}(\pi_{1}^{!}F,\pi_{2}^{*}G)$
@@
4.3. Microsheaves on polarized real contact manifolds
A key tool of [15] is the functoriality of $\mu sh$ under (quantized) contact transformation [15, Sec. 7]. One readily obtains:
@@ 4.3. Microsheaves on polarized real contact manifolds
Suppose given a contact manifold $V$ and a contact embedding $\iota:V\times T^{*}D^{n}$, for some disk $D$. Then $\mu sh_{\iota(V)\times D^{n}}$ is locally constant along $D^{n}$, hence the pullback of a sheaf of categories along $V\times D^{n}\to V$.
The basic idea of [26] is to use Lemma 4.9 as a definition of $\mu sh_{V}$. Let $(V,\xi)$ be a contact manifold. Suppose given a (possibly positive codimensional) contact embedding $\iota:V\to S^{*}M$, and any Lagrangian distribution $\eta$ of the symplectic normal bundle to $\iota$. Let us write $V(\eta)$ for a thickening of $\iota(V)$ along $\eta$ in a tubular neighborhood. Then $\iota^{*}\mu sh_{V(\eta)}$ provides a sheaf of categories on $V$. By contact transformations, it is easy to see that this sheaf of categories is locally equivalent to microsheaves on a contact cosphere bundle of the same dimension.
@@
To eliminate the dependence on $\iota$ and $M$, one notes that Gromov’s $h$-principle for contact embeddings implies the existence of high codimension embeddings of $V$ into the standard contact $\mathbb{R}^{2n+1}$ for large enough $n$, the space of which moreover becomes arbitrarily connected as $n\to\infty$. There remains the (continuous) dependence on a choice of polarization of the stable symplectic normal bundle (“stable normal polarization”).
@@
Above, $-\eta$ can be regarded as just a notational choice. To give it an actual meaning, note that $BO\to BU$ is a morphism of spectra, so given a stable symplectic vector bundle $E$ on some topological space $X$, say classified by some map $E:X\to BU$, and a polarization of $E$, i.e. lift to some $F:X\to BO$, then $-F$ gives a polarization of $-E$, where -- is the pullback by the canonical ‘inverse’ involution on $BO$ or $BU$.
In fact [22, Sec. 10.2], it is always possible to build a contact manifold, $V^{T^{*}LGr(\xi)}$, which is a bundle over $V$ with fibers the cotangent bundles to the Lagrangian Grassmannians of the contact distribution $\xi$. The fiberwise zero section is the Lagrangian Grassmannian bundle $V^{LGr(\xi)}$. The virtue of $V^{T^{*}LGr(\xi)}$ is that it has a canonical polarization of the contact distribution, hence a canonical stable normal polarization. Thus we define $\mu sh_{V^{T^{*}LGr(\xi)}}$, and, restricting supports, $\mu sh_{V^{LGr(\xi)}}$.
Let us explain how this recovers the previous notion. A polarization $\rho$ of the contact distribution is (by definition) a section $\rho:V\to V^{LGr(\xi)}$. Now, the normal bundle to $\rho(V)\subset V^{T^{*}LGr(\xi)}$ is $T^{*}LGr(\xi)$, this has the canonical polarization by the cotangent fiber. We may combine this with canonical polarization of the symplectic stable normal bundle of $V^{T^{*}LGr(\xi)}$ to obtain a polarization of the symplectic stable normal bundle of $V$. It’s an exercise to see that this polarization is canonically identified with $-\rho$. We conclude:
@@
We may choose the embedding of $V$ by first embedding $V^{LGr(\xi)}$. To obtain $\mu sh_{V,\rho}$, we must then thicken $V$ along a polarization of its normal bundle which stabilizes to $-\rho$. We obtain such by thickening $V$ along $V^{LGr(\xi)}\subset V^{T^{*}LGr(\xi)}$ and then along the canonical polarization of the normal bundle to $V^{T^{*}LGr(\xi)}$. But the same procedure defines $\mu sh_{V^{LGr(\xi)}}$. ∎
@@
Let $M$ be a manifold, and $\nu$ the stable normal bundle of $M$. Then the stable symplectic normal bundle to $S^{*}M$ is $\nu\oplus\nu^{*}$, and the stable normal polarization by $\nu$ is canonically identified with $-\phi$.
Corollary 4.13.
There is a canonical equivalence of categories $\mu sh_{S^{*}M}\cong\mu sh_{S^{*}M,\phi}$, where the LHS is defined by Equation 9 and the right hand side by Theorem 4.10.

### v3

@@ 4.1. Sheaves on manifolds
Let $M$ be a real manifold. Fix a symmetric monoidal stable presentable compactly generated category, $\mathcal{C}$. The reader will not lose much of the point of the paper taking throughout $\mathcal{C}$ to be the derived category of dg modules over some commutative ring $R$. We write $sh(M)$ for the (stable) category of sheaves on $M$ with values in $\mathcal{C}$.
@@ 4.1.1. Microsupport
The microsupport is easily seen to be conical and satisfy $ss(Cone(F\to G))\subset ss(F)\cup ss(G)$. A deep result of [25, Thm. 6.5.4] is that the microsupport is coisotropic (also called involutive; see [25, Def. 6.5.1]). For a conic subset $K\subset T^{*}M$, we write $sh_{K}(M)$ for the full subcategory on objects microsupported in $K$. For any subset $\Lambda\subset S^{*}M$, we write $sh_{\Lambda}(M):=sh_{\mathbb{R}_{+}\Lambda\cup 0_{M}}(M)$, with $0_{M}$ the zero section of the cotangent bundle. We write $T^{\circ}M:=T^{*}M-0_{M}$; following the usual convention, we will not distinguish between subsets of $S^{*}M$ and conic subsets of $T^{*}M-0_{M}$.
@@ 4.2. Microsheaves on cotangent bundles
Let $\mu sh_{T^{*}M}$ be the sheaf of categories on $T^{*}M$ defined by sheafifying the presheaf $sh_{T^{*}M}^{pre}(\Omega)$ in (9). Similarly, let $\mu sh_{S^{*}M}$ be the presheaf of categories on $S^{*}M$ obtained by sheafifying $\mu sh^{pre}_{S^{*}M}$.
In any sheaf of categories $\mathcal{X}$ on a topological space $T$, given $F,G\in\mathcal{X}(T)$, the assignment $U\mapsto\operatorname{Hom}_{\mathcal{X}(U)}(F|_{\Omega},G|_{\Omega})$ is a sheaf on $T$; let us denote it as $\mathcal{H}om_{\mathcal{X}}(F,G)$. A fundamental result of Kashiwara and Schapira computes this Hom sheaf via sheaf operations [25, Thm. 6.1.2]: [^1] More precisely, [25] shows there is a morphism of this kind for $\mu sh^{pre}$ which is an isomorphism at stalks; the stated result follows upon sheafification. See [35] for some detailed discussions about the sheafification of $\mu sh^{pre}$.
(10) $\mathcal{H}om_{\mu sh}(F,G)=\mu hom(F,G):=\mu_{\Delta}\operatorname{Hom}_{M\times M}(\pi_{1}^{*}F,\pi_{2}^{!}G).$
@@
The sheaf $\mu sh_{T^{*}M}$ is conic, i.e. equivariant for the $\mathbb{R}^{+}$ scaling action. In particular, $\mu sh_{T^{*}M}|_{T^{\circ}M}$ is locally constant in the radial direction. This being a contractible $\mathbb{R}^{+}$, we may define a sheaf of categories $\mu sh_{S^{*}M}$ on $S^{*}M$ equivalently by pushforward or pullback along an arbitrary section of $T^{\circ}M\to S^{*}M$. As $ss(G)$ is conic for $G\in\mu sh_{S^{*}M}$, it uniquely determine a set $ss^{\infty}(G)\subseteq S^{*}M$.
@@
[25, Prop. 6.6.1] Let $M$ be a manifold and let $N\subset M$ be a submanifold. Let $L=T^{*}_{N}M$ and fix a point $p\in L$. Then there is an equivalence of categories:
(11) $\mathcal{C}$ $\xrightarrow{\sim}(\mu sh_{L})_{p}$
@@
Alternatively, one can also restrict to constructible sheaves all from the beginning and define
(12) $\mu sh_{T^{*}M,\mathbb{R}-c}^{pre}(\Omega):=sh_{\mathbb{R}-c}(M)/sh_{T^{*}M\setminus\Omega,\mathbb{R}-c}(M),$
and consider its sheafification. For a conic Lagrangian $\Lambda\subseteq T^{*}M$, define
(13) $\mu sh_{\mathbb{R}-c,\Lambda}^{pre}(\Omega):=sh_{\mathbb{R}-c,\Lambda\cup(T^{*}M\setminus\Omega)}(M)/sh_{\mathbb{R}-c,T^{*}M\setminus\Omega}(M).$
Unwrapping the definition, we see that $sh_{\mathbb{R}-c,\Lambda\cup(T^{*}M\setminus\Omega)}(M)=\{F\in sh_{\mathbb{R}-c}(M)|ss(F)\cap\Omega\subseteq\Lambda\}$. Similarly, one can define $\mu sh^{pre}_{T^{*}M,\mathbb{C}-c}$ when $M$ is complex, and $\mu sh^{pre}_{\Lambda,\mathbb{C}-c}$ when $\Lambda$ is complex. By microlocal cut-off, these notions agree with the ones given previously.
@@
[42, Theorem 3.2.2] The canonical map $\mu sh_{T^{*}M,\mathbb{R}-c}^{pre}\rightarrow\mu sh_{T^{*}M,\mathbb{R}-c}$ induces an isomorphism upon sheafification. The similar statement holds for $\mu sh_{\mathbb{R}-c,\Lambda}$, $\mu sh_{T^{*}M,\mathbb{C}-c}$, and $\mu sh_{\mathbb{C}-c,\Lambda}$.
We will need the following lemma in the next section. Let $i:N\hookrightarrow M$ be an inclusion of closed submanifolds. The map $i$ induces an inclusion $i_{\pi}:T^{*}M|_{N}\hookrightarrow T^{*}M$ and the transpose of its derivative induces a projection
$(di)^{t}:T^{*}M|_{N}$ $\rightarrow T^{*}N$
$(y,\xi)$ $\mapsto(y,\xi\circ di_{y}).$
Lemma 4.9.
The *-pushforward, $i_{*}:sh(N)\xrightarrow{\sim}sh_{N}(M)$, microlocalizes to an equivalence
$(di)^{t}]^{*}\mu sh_{T^{*}N}\xrightarrow{\sim}\mu sh_{T^{*}M|_{N}}.$
(Here $\mu sh_{T^{*}M|_{N}}\subset\mu sh_{T^{*}M}$ is the subsheaf of full subcategories on objects supported in $T^{*}M|_{N}\subset T^{*}M$; see Definition 4.5.)
Proof.
Since $i$ is a closed embedding, [25, Proposition 5.4.4] implies that $ss(i_{*}F)=({(di)^{t}})^{-1}ss(F)$ is contained in $T^{*}M|_{N}$ for $F\in sh(N)$. This implies that the assignment
$\mu sh_{T^{*}N}^{pre}(\Omega)$ $\rightarrow\left(({(di)^{t}})_{*}{i_{\pi}}^{*}\mu sh_{T^{*}M|_{N}}^{pre}\right)(\Omega)$
$F$ $\mapsto({i_{\pi}})_{*}F$
is well-defined. To check the induced map $({(di)^{t}})^{*}\mu sh_{T^{*}N}\rightarrow{i_{\pi}}^{*}\mu sh_{T^{*}M}$ on sheaves is an equivalence, one can check at stalks. Fully faithfulness is then implied by [25, Proposition 4.4.7(ii)], as (10) shows that the Hom is computed by $\mu hom$. Essential surjectivity is implied by [25, Prop. 6.6.1]. ∎
Before we leave this section, we mention two common tools for studying microsheaves. The first is the contact transformation.
Theorem 4.10 ([25, Corollary 7.2.2]).
Let $\mathcal{U}\subseteq S^{*}M$, and $\mathcal{V}\subseteq S^{*}N$ be open sets, and $\chi:\mathcal{U}\xrightarrow{\mathcal{V}}$ be a contactomorphism. Then, for any given $p\in\mathcal{U}$, shrink $\mathcal{U}$ if needed, one can assume that there exists a sheaf $K\in sh(M\times N)$ such that the functor $\Phi_{K}:sh(M)\rightarrow sh(N)$ given by convolving with $K$ induces an equivalence
$\Phi_{K}:\mu sh^{pre}_{S^{*}M}|_{\mathcal{U}}\xrightarrow{\sim}\chi^{*}\left(\mu sh^{pre}_{S^{*}N}|_{\mathcal{V}}\right).$
Consequently, it induces an equivalence $\mu sh_{S^{*}M}|_{\mathcal{U}}\xrightarrow{\sim}\chi^{*}\mu sh_{S^{*}N}|_{\mathcal{V}}$ which commutes with the canonical map $\mu sh^{pre}\rightarrow\mu sh$.
The other one is the doubling trick, which states that, if $\mathcal{F}\in\mu sh(\Omega)$, for some open set $\Omega\subseteq S^{*}M$, is a microsheaf whose support $ss^{\infty}(\mathcal{F})$ is contained in some closed Legendrian $\Lambda\subseteq S^{*}M$, then $\mathcal{F}$ can be represented by a genuine sheaf $F$ (with possibly larger microsupport).
Theorem 4.11 ([35, Theorem 7.18], [28, Theorem 4.47] ).
Let $\Lambda\subseteq S^{*}M$ be a closed Legendrian and $\Omega\subseteq S^{*}M$ an open set. Choose a positive isotopy $\Lambda_{t}$ of $\Lambda$ in $S^{*}M$ such that $\alpha(\partial_{t}\Lambda_{t})$ is nonzero on $\Omega$ and outside $\Omega$. Then, for some small enough $\epsilon$, the composition
$sh_{\Lambda\cup\Lambda_{\epsilon}}(M)\rightarrow\mu sh_{\Lambda\cup\Lambda_{\epsilon}}(\Omega)\rightarrow\mu sh_{\Lambda}(\Omega)$
is surjective. Here, we use the fact that $\Lambda$ and $\Lambda_{\epsilon}$ are disjoint in $\Omega$ for the second arrow.
4.3. Microsheaves and Maslov index
In this subsection, we describe following [6] the sheaf quantization of a rotation of $J^{1}\mathbb{R}^{n}$. This material will be used explicitly only in Section 5.2.
Endow $T^{*}\mathbb{R}^{n}\times T^{*}\mathbb{R}$ with coordinates $(x,\xi,t,\tau)$. Identify the $1$-jet space $J^{1}\mathbb{R}^{n}$ with the hyperplane $\{\tau=1\}\subset T^{*}\mathbb{R}^{n}\times T^{*}\mathbb{R}$; the contact form restricted to $\{\tau=1\}$ is
$\alpha\;=\;\xi\cdot dx+dt.$
Fix $1\leq\ell\leq n$. As in [6], we consider a $1$-parameter family of contactomorphisms $\psi_{t}:J^{1}\mathbb{R}^{n}\to J^{1}\mathbb{R}^{n}$ $\psi_{t}:J^{1}\mathbb{R}^{n}\to J^{1}\mathbb{R}^{n}$ given by
$\psi_{t}(x,\xi,\tau)\;=\;(x^{\prime},\xi^{\prime},\tau^{\prime})$
with (for $i=1,\dots,\ell$)
$x^{\prime}_{i}$ $=(\cos t)\,x_{i}-(\sin t)\,\xi_{i},$
$\xi^{\prime}_{i}$ $=(\sin t)\,x_{i}+(\cos t)\,\xi_{i},$
and for $i=\ell+1,\dots,n$
$x^{\prime}_{i}=x_{i},\qquad\xi^{\prime}_{i}=\xi_{i}.$
The $\tau$-coordinate transforms by
$\tau^{\prime}\;=\;\tau+F_{t}(x,\xi),$
where,
(14) $F_{t}(x,\xi)=\sum_{i=1}^{\ell}\left(\frac{\sin t\cos t}{2}\,(x_{i}^{2}-\xi_{i}^{2})-2\sin^{2}t\cdot x_{i}\cdot\xi_{i}\right).$
Note that $\psi_{t}$ is the lift of a $1$-parameter family of sympectomorphisms of $T^{*}\mathbb{R}^{n}$ which rotates the $(x_{i},\xi_{i})$ coordinates clockwise for $1\leq i\leq\ell$.
These symplectomorphisms admit a sheaf quantization:
Proposition 4.12 ([6, Proposition 3.10] [^2] As the Hamiltonian is not compactly supported, the quantization does not follow directly from the usual result of Guillermou, Kashiwara, and Schapira [15, Proposition A.6].).
Consider the space $\mathbb{R}^{n}\times\mathbb{R}^{n}\times\mathbb{R}_{t}\times\mathbb{R}_{s}$ where we view $s$ as the time direction. Then, there exists a sheaf kernel $\mathcal{S}\in sh(\mathbb{R}^{n}\times\mathbb{R}^{n}\times\mathbb{R}_{t}\times\mathbb{R}_{s})$ such that
  1. (1)
    $\mathcal{S}|_{s=0}=1_{\Delta_{\mathbb{R}^{n}}}\boxtimes 1_{[0,\infty)}$.
  2. (2)
    $ss(\mathcal{S})\cap\{\sigma>0\}\subseteq\Lambda_{\Psi}$ where we use $\sigma$ to denote the cotangent coordinate for $\mathbb{R}_{s}$ and $\Lambda_{\Psi}$ is the Legendrian lift of the isotopy movie of the Hamiltonian rotation.
  3. (3)
    Over the open interval $s\in(0,\pi)$, we have the explicit formula,
    $\mathcal{S}|_{(0,\pi)_{s}}=1_{\{(x,y,t,s)|t+F_{2s}(x,y)\geq 0\}}.$
For a sheaf kernel $\mathcal{K}\in sh(M\times N\times\mathbb{R}_{t})$, one can consider the $\star$-convolution
$sh(M\times\mathbb{R}_{t_{1}})$ $\rightarrow sh(N\times\mathbb{R}_{t})$
$\mathcal{F}$ $\mapsto\mathcal{K}\star\mathcal{F}\coloneqq{(p_{N,+})}_{!}(p_{2}^{*}\mathcal{K}\otimes p_{M,1}^{*}\mathcal{F})$
where $p_{M,1}$ and $p_{2}$ are the usual projections from $M\times N\times\mathbb{R}_{t_{1}}\times\mathbb{R}_{t_{2}}$ to $M\times\mathbb{R}_{t_{1}}$ and $M\times N\times\mathbb{R}_{t_{2}}$ but, when pushing forward, we use the addition for the $t$ direction by $p_{N,+}(x,y,t_{1},t_{2})=(y,t)$, $t=t_{1}+t_{2}$ instead.
By (2) of the above Proposition 4.12, $\mathcal{S}\star(-)$ restricts to an equivalence on $\mu sh^{pre}|_{\{\sigma>0\}}$ and hence to the restriction of $\mu sh$ to $\{\sigma>0\}$.
Let $K:=\{\xi=0,\ \tau=0\}=0_{\mathbb{R}^{n}}\times\{0\}\subset J^{1}\mathbb{R}^{n}$ be the zero section, and write $p=(0,0,0)$ for the origin. Set $K_{t}:=\psi_{t}(K)$ and note that $K_{\pi/2}$ is the conormal to $C_{k}\coloneqq\{x_{1}=\cdots=x_{k}=0\}\subset\mathbb{R}^{n}$.
Lemma 4.13.
We have the following diagram:
(15) $\mathcal{C}$$(\mu sh_{K})_{p}$$\mathcal{C}$$(\mu sh_{K_{\pi/2}})_{p}$$\scriptstyle{A\mapsto A_{L}\boxtimes 1_{[0,\infty)}}$$\scriptstyle{[-k]}$$\scriptstyle{A\mapsto A_{C_{k}}\boxtimes 1_{[0,\infty)}}$
where the right downward arrow is induced by the sheaf quantization of $\psi_{\pi/2}$ furnished by Proposition 4.12.
Proof.
By the Künneth formula, it is sufficient to prove the case for $l=1$. Since we are rotating the Lagrangian by $90^{o}$, we set $s=\pi/2$ and we have $\mathcal{S}|_{s=\pi/2}=1_{\{t-\frac{1}{2}x\xi\}}$ by (14). Thus, seeing the commutativity amounts to computing the object
$\mathcal{S}|_{s=\pi/2}\star(A_{\mathbb{R}^{1}}\boxtimes 1_{[0,\infty)_{t}})=(p_{2,+})_{!}\left(A_{\{(x,y,t_{1},t_{2})|t_{1}\geq 0,\ t_{2}-\frac{1}{2}\geq 0\}}\right).$
Now, the projection $p_{2,+}$ can be decomposed to the naive projection $q(x,y,t_{1},t_{2})=(y,t_{1},t_{2})$ follows by the addition $a(y,t_{1},t_{2})=(y,t_{1}+t_{2})$, we can thus compute the pushforward in two steps. The claim is that $q_{!}\left(A_{\{(x,y,t_{1},t_{2})|t_{1}\geq 0,\ t_{2}-\frac{1}{2}xy\geq 0\}}\right)=A_{\{0\}}[-1]\boxtimes 1_{[0,\infty)_{t_{1}}}\boxtimes 1_{[0,\infty)_{t_{2}}}$, and we will be done because the projection formula will imply that
$a_{!}\left(A_{\{0\}}[-1]\boxtimes 1_{[0,\infty)_{t_{1}}}\boxtimes 1_{[0,\infty)_{t_{2}}}\right)=A_{\{0\}}[-1]\boxtimes(1_{[0,\infty)_{t_{1}}}\star 1_{[0,\infty)_{t_{2}})})=A_{\{0\}}[-1]\boxtimes 1_{[0,\infty)_{t}}$
as we desired. To see the claim, again by the projection formula, we can ignore $t_{1}$ and we only have to show that, if $f(x,y,t)\coloneqq(y,t)$, then $f_{!}(A_{\{t-\frac{1}{2}xy\geq 0\}})=A_{(\{0,t)|t\geq 0\}}[-1]=A_{\{0\}}[-1]\boxtimes 1_{[0,\infty)}$. Lastly, we can check on stalks, and consider, for $(y,t)\in\mathbb{R}^{2}_{(y,t)}$,
$\left(f_{!}(A_{\{t-\frac{1}{2}xy\geq 0\}})\right)_{(y,t)}=\Gamma_{c}(\{x\in\mathbb{R}|t-(1/2)xy\geq 0\};A)$
the compactly supported sections on the set $\{x\in\mathbb{R}|t-\frac{1}{2}xy\geq 0\}$. We note that, when $y>0$, $t-\frac{1}{2}xy$ is equivalent to $x\leq\frac{2t}{y}$ so for any pair $(y,t)$ the set $\{x\in\mathbb{R}|t-\frac{1}{2}xy\geq 0\}$ is topologically $0,\infty)$ which has a vanishing compactly supported cohomology. The same situation holds when $y<0$. Thus, the only non-trivial stalks live over $y=0$, in which case $x$ has no condition. In other words, the stalk is given by $\Gamma_{c}(\mathbb{R};A)=A[-1]$ and is exactly what we wanted. ∎
Remark 4.14.
Lemma 4.13 can presumably also be extracted from the discussion in [25, Sec. 7.5, Appendix A]. Similar computations have also appeared in e.g. [14, 17].
4.4. Microsheaves on polarized (real) contact manifolds
In the previous section, we discussed microlocal sheaves on cotangent/cosphere bundles. The high-codimensional embedding trick of [39] extends the definition of microlocal sheaves to arbitrary contact or exact symplectic manifolds equipped with Maslov data. In this section, we review how this works on polarized contact manifolds; we postpone the discussion of Maslov data to the next section.
The starting point of [39] is the following lemma, which is a consequence of the functoriality of $\mu sh$ under (quantized) contact transformation [25, Sec. 7].
Lemma 4.15.
@@ 4.4. Microsheaves on polarized (real) contact manifolds
Suppose given a contact manifold $V$ and a contact embedding $\iota:V\times T^{*}D^{n}\hookrightarrow S^{*}M$, for some disk $D^{n}$ and some manifold $M$. Then $\mu sh_{\iota(V\times D^{n})}$ is locally constant along $D^{n}$, hence the pullback of a sheaf of categories along $V\times D^{n}\to V$.
Proof.
The statement is local, so we may assume $V$ is a Darboux ball. By the contact neighborhood theorem [13, Theorem 2.5.15], there is an open ball $U\subset S^{*}\mathbb{R}^{n+m}$ and a contactomorphisms $(V\times T^{*}D^{n},V\times\{0\})\simeq(U,U\cap S^{*}\mathbb{R}^{n})$, where $S^{*}\mathbb{R}^{m}\subset S^{*}\mathbb{R}^{n+m}$ is induced by the standard inclusion $\mathbb{R}^{m}=\mathbb{R}^{n}\times\{0\}^{n}\hookrightarrow\mathbb{R}^{m+n}$. By the functoriality of $\mu sh$ under (quantized) contact transformation [25, Sec. 7], we may therefore assume without loss of generality that $V=U\cap S^{*}\mathbb{R}^{n}$. But, by Lemma 4.9, we conclude that $\mu sh_{S^{*}\mathbb{R}^{m}\times 0_{\mathbb{R}^{n}}}=q^{*}\mu sh_{S^{*}\mathbb{R}^{m}}$ where $q:S^{*}\mathbb{R}^{m}\times T^{*}\mathbb{R}^{n}\rightarrow S^{*}\mathbb{R}^{m}$ is the projection. ∎
The basic idea of [39] is to use Lemma 4.15 as a definition of $\mu sh_{V}$. Let $(V,\xi)$ be a contact manifold. Suppose given a (possibly positive codimensional) contact embedding $\iota:V\to S^{*}M$ and any Lagrangian distribution $\eta$ of the symplectic normal bundle $N_{V}$ to $\iota$. The choice of $\eta$ comes with a coisotropic embedding $\tau_{\eta}:\operatorname{Op}(0_{V}\cap\eta)\to S^{*}M$ such that the differential satisfies $\operatorname{im}(d\tau_{\eta}:\{0\}\times\eta(p)\subset T_{p}V\times(N_{V})_{p}\to T_{\tau_{\eta}(p,0)}S^{*}M)=\eta(p),$ for all $p\in V$. Here $0_{V}$ is the zero section of the symplectic normal bundle $N_{V}$. We denote by $V(\eta)$ the image of $\tau_{\eta}$ and call it a thickening of $\iota(V)$ along $\eta$.
Lemma 4.16.
The sheaf of categories $\mu sh_{V(\eta)}$ is the pullback of certain a sheaf of categories on $V$. We will denote this sheaf as $\mu sh_{V,\iota,\eta}$.
Proof.
The Lagrangian bundle structure $\eta$ provides, near $V$, a contraction $r:V(\eta)\rightarrow V$, and thus defines a sheaf $\mu sh_{V,\iota,\eta}\coloneqq r_{*}\mu sh_{V(\eta)}$ by pushforward. However, the map $r$ is locally given by the projection $V\times D^{n}\rightarrow V$, and we’ve shown in Lemma 4.15 that the sheaf $\mu sh_{V(\eta)}$ is constant along the fiber direction, so $\mu sh_{V(\eta)}$ can be recovered from $\mu sh_{V,\iota,\eta}$ by
$\mu sh_{V(\eta)}=r^{*}r_{*}\mu sh_{V(\eta)}=r^{*}\mu sh_{V,\iota,\eta}.$
This completes the proof. ∎
@@
To eliminate the dependence on $\iota$ and $M$, one notes that Gromov’s $h$-principle for contact embeddings implies the existence of high codimension embeddings of $V$ into the standard contact $\mathbb{R}^{2n+1}$ for large enough $n$, the space of which moreover becomes arbitrarily connected as $n\to\infty$. There remains the (continuous) dependence on a choice of polarization of the stable symplectic normal bundle (“stable normal polarization”). That is:
@@
Above, the sign - on $\eta$ can be regarded as just a notational choice. To give it an actual meaning, note that $BO\to BU$ is a morphism of spectra, so given a stable symplectic vector bundle $E$ on some topological space $X$, say classified by some map $E:X\to BU$, and a polarization of $E$, i.e. lift to some $F:X\to BO$, then F-F gives a polarization of E-E, where - is the pullback by the canonical ‘inverse’ involution on $BO$ or $BU$.
In fact [35, Sec. 10.2], it is always possible to build a contact manifold, $V^{T^{*}LGr(\xi)}$, which is a bundle over $V$ with fibers the cotangent bundles to the Lagrangian Grassmannians of the contact distribution $\xi$. The fiberwise zero section is the Lagrangian Grassmannian bundle $V^{LGr(\xi)}$. The virtue of $V^{T^{*}LGr(\xi)}$ is that it has a canonical polarization of the contact distribution, hence a canonical stable normal polarization. Thus we define $\mu sh_{V^{T^{*}LGr(\xi)}}$, and, restricting supports, $\mu sh_{V^{LGr(\xi)}}$. It is evident from the construction that $\mu sh_{V^{LGr(\xi)}}$ is locally constant in the Lagrangian Grassmannian direction. We may also stabilize $\xi\to\xi\oplus T^{*}\mathbb{R}^{n}$; taking $n\to\infty$, we have a sheaf of categories $\mu sh_{V^{LGr}}$ on the (stable) Lagrangian Grassmannian bundle $V^{LGr}$, locally constant along the Lagrangian Grassmannian direction.
Let us explain how this recovers the previous notion. A polarization $\rho$ of the contact distribution is (by definition) a section $\rho:V\to V^{LGr(\xi)}$. Now, the normal direction of a neighborhood to $\rho(V)\subset V^{T^{*}LGr(\xi)}$ can be identified with $T^{*}LGr(\xi)$. This carries the canonical polarization by the cotangent fiber. We may combine this with canonical polarization of the symplectic stable normal bundle of $V^{T^{*}LGr(\xi)}$ to obtain a polarization of the symplectic stable normal bundle of $V$. It’s an exercise to see that this polarization is canonically identified with $\rho$. We conclude:
@@
We may choose the embedding of $V$ by first embedding $V^{LGr(\xi)}$. To obtain $\mu sh_{V,\rho}$, we must then thicken $V$ along a polarization of its normal bundle which stabilizes to $\rho$. We obtain such by thickening $V$ along $V^{LGr(\xi)}\subset V^{T^{*}LGr(\xi)}$ and then along the canonical polarization of the normal bundle to $V^{T^{*}LGr(\xi)}$. But the same procedure defines $\mu sh_{V^{LGr(\xi)}}$. ∎
@@
Let $M$ be a manifold, and $\nu$ the stable normal bundle of $M$. Then the stable symplectic normal bundle to $S^{*}M$ is $\nu\oplus\nu^{*}$, and the stable normal polarization by $\nu$ is canonically identified with $\phi$.
Proposition 4.20.
There is a canonical equivalence of categories $\mu sh_{S^{*}M}\simeq\mu sh_{S^{*}M,\phi}$, where the LHS is defined by (9) and the right hand side by Theorem 4.17.